OpenStudy (anonymous):

what volume of oxygen is needed to react with 40cm3 of methane and what volume of CO2 is formed if all gasses are measured up at room temp and pressure

OpenStudy (cherylim23):

Recognise the combustion equation: CH4 (g) + 2 O2 (g) --> CO2 (g) + 2 H2O (g) n(O2) = n(CH4) x 2 This implies that v(O2) = v(CH4) x 2 = 40 x 2 = 80cm3 since all products and reactants are at gaseous state at rtp

OpenStudy (anonymous):

i dont get it can you please explain in terms of molar eq plz

OpenStudy (cherylim23):

Because all reactants and products are at gaseous state, you can assume that their volumes are directly proportionate to their amount in moles, since moles = volume/molar volume. Does this make sense?

OpenStudy (anonymous):

yes thank u :D